OpenStudy (anonymous):

Find the vertex of the parabola. y=x²-6x+7

OpenStudy (anonymous):

SO this is the standard form for parabolas f(x) = ax² + bx + c

OpenStudy (anonymous):

V=(-b)/(2a) put the two together bud

OpenStudy (anonymous):

y=x²-6x+7 a= 1 b= -6 vertex: x = -b/2a x= -(-6) / 2(1) x=3

OpenStudy (anonymous):

now plau that in for x in the original equation

OpenStudy (anonymous):

*plug

OpenStudy (anonymous):

so y= 3^2(squarred) - 6 (3) + 7 y= -2 So vertex = (3,-2)